# Graph Reinforcement Learning for Combinatorial Optimization: A Survey and Unifying Perspective

### v1

@@ 5.1 Challenges of Graph RL
5.1.1 Framing Problems as MDPs
In order to apply RL for a given problem, one needs to decide how to frame it as a Markov Decision Process, which will impact the learning effectiveness. While the objective function is typically dictated by the application, the designer generally has the liberty to decide the state and action spaces as well as the dynamics. Let us discuss some general considerations.
@@ 5.1.1 Framing Problems as MDPs
To reduce the action space dimension, other works in deep RL have considered generalizing across similar actions by embedding them in a continuous space (Dulac-Arnold et al., 2015) and learning which actions should be eliminated via supervision provided by the environment (Zahavy et al., 2018). Existing approaches in planning consider progressively widening the search based on a heuristic (Chaslot et al., 2008) or devising a policy for eliminating actions in the tree (Pinto & Fern, 2017; Darvariu et al., 2023a).
5.1.2 Reward Design
@@
5.1.3 Choosing and Designing RL Algorithms and Learning Representations
Given an MDP formulation, how does one choose an RL algorithm to solve it? Since the application of RL to graph combinatorial optimization problems is still a nascent area of investigation, this complex question is often glossed over by papers. The novelty of casting problems as MDPs and using RL is often deemed a sufficient contribution, especially as this methodology can already yield gains in evaluation metrics compared to classic techniques. However, the literature as a whole lacks systematic comparisons of RL techniques for combinatorial optimization problems, or guidelines on which algorithm to pick depending on the characteristics of the problem.
@@ 5.1.3 Choosing and Designing RL Algorithms and Learning Representations
    Among model-free algorithms, off-policy algorithms (e.g., DQN) are more sample-efficient than on-policy algorithms (e.g., PPO), meaning that they will require less environment interactions to reach a well-performing policy. If evaluating the objective function is computationally expensive, as is often the case for Graph RL problems, then using an off-policy algorithm is advisable.
@@
Symmetries that are present in the problem and solution spaces are also highly relevant and can be fruitfully exploited by algorithms and learning representations alike. For example, a TSP solution that visits nodes $(A,B,C)$ is equivalent to one that visits $(B,C,A)$. Exploiting this may lead to better sample complexity and more robust models. Approaches for canonical problems have considered encoding symmetry with specific terms in the loss function (Kwon et al., 2020; Kim et al., 2022) or augmenting the dataset of sampled solutions at inference time (Kwon et al., 2020). A recent work (Drakulic et al., 2023) proposed exploiting symmetries at the level of the MDP formulation itself by proposing a transformation of the original MDP that reduces the state space. Symmetries may also be encoded in the learning representation itself, as can be achieved by the use of GNNs with appropriate permutation-invariant readout functions (Khalil et al., 2017).
5.1.4 Scalability
@@
Let us enumerate some further possibilities beyond those outlined in the challenges above. One can study the application of models trained on small problem instances to larger ones. This technique requires models that are independent of the size of the problem, and can obtain impressive results in scenarios in which the model transfers well, but can equally suffer from substantial degradation in solution quality if the structure of the solution space is dissimilar at varying scales. Further options that can be adopted are the use of decision-time planning to examine a small fraction of the entire decision-making process, and using demonstrations of a well-performing algorithm to collect data instead of online environment interaction.
5.1.5 Generalization
An important issue is the fact that one cannot guarantee that the learned models will generalize well when encountering instances outside of the training distribution. This fundamental limitation is reminiscent of the No Free Lunch theorems (Wolpert & Macready, 1997) in Supervised Learning, which suggest that it is not possible to obtain a model that performs well in all possible scenarios. Partial mitigation may be possible by training models on a wide variety of scenarios including different initial starting points for the RL agent in the environment (i.e., different initial solutions), problem instances of varying difficulty and size, and even different variants of related problems. However, this fundamental limitation remains, as we cannot train a model for a scenario we know nothing about. In settings where such models are deployed, one may want to have the tools in place to detect distribution shift, and possibly use a fall-back approach whose properties and expected performance are well-understood.
@@ 6 Conclusion and Outlook
When are such approaches useful? An overall picture is that such methods are advantageous for problems with no known powerful exact algorithms, and for which heuristic or metaheuristics are typically used in the literature. These tend to be problems with objectives and constraints that are not “well-behaved” mathematically speaking, e.g., problems with nonconvex and nonlinear facets. These aspects make them difficult to solve analytically, but are possibly amenable to be approximated using deep neural networks.
Adopting a critical stance, it is difficult to know a priori whether a given problem is amenable to an RL-based solution, or how much of a gain we might obtain over a classic approach. Given their technical complexity and experimental challenges, RL methods may require several development cycles until satisfactory performance is obtained. Considering the current popularity of RL, one might also ask if their application in this problem space is an instance of Maslow’s hammer (Maslow, 1966). Is it the case that we favor the ubiquitous use of this tool over careful appreciation of what might be the appropriate methodology for a given set of problems? Nevertheless, the potential of RL approaches in this problem space is significant and transformative, a belief that is strongly supported by the emerging body of literature. As RL techniques become more widespread, we expect them to find successful applications far beyond canonical problems, and to transform the scientific discovery process (Wang et al., 2023).
Can we, therefore, collectively hang up our boots, leaving the machines to discover how to solve these problems? We argue that this not the case. Generic decision-making algorithms and learning representations are clearly not a silver bullet since they do not necessarily exploit problem structure efficiently. There are substantial improvements to be made by encoding knowledge and understanding about the problem into these solution approaches. We envisage the future of algorithm development as leveraging both deep problem insights and highly efficient machine-learned components.

### v2

@@ 5.1 Challenges of Graph RL
5.1.1 Framing Graph Combinatorial Optimization Problems as MDPs
In order to apply RL for a given graph combinatorial optimization problem, one needs to decide how to frame it as a Markov Decision Process, which will impact the learning effectiveness. While the objective function is typically dictated by the application, the designer generally has the liberty to decide the state and action spaces as well as the dynamics. Let us discuss some general considerations.
@@ 5.1.1 Framing Graph Combinatorial Optimization Problems as MDPs
To reduce the action space dimension, other works in deep RL have considered generalizing across similar actions by embedding them in a continuous space (Dulac-Arnold et al., 2015) and learning which actions should be eliminated via supervision provided by the environment (Zahavy et al., 2018). Existing approaches in planning consider progressively widening the search based on a heuristic (Chaslot et al., 2008) or devising a policy for eliminating actions in the tree (Pinto & Fern, 2017). Such techniques have also been applied for Graph RL (Darvariu et al., 2023a), but remain largely unexplored.
5.1.2 Reward Design: Balancing Accuracy, Speed, and Multiple Objective Functions
@@
5.1.3 Choosing and Designing Algorithms and Learning Representations for Graph RL
Given an MDP formulation, how does one choose an RL algorithm to solve it? Since the application of RL to graph combinatorial optimization problems is still a nascent area of investigation, this complex question is often glossed over by papers. The literature as a whole lacks systematic comparisons of RL techniques for combinatorial optimization problems, or guidelines on which algorithm to pick depending on the characteristics of the problem.
@@ 5.1.3 Choosing and Designing Algorithms and Learning Representations for Graph RL
    Among model-free algorithms, off-policy algorithms (e.g., DQN) are more sample-efficient than on-policy algorithms (e.g., PPO), meaning that they will require fewer environment interactions to reach a well-performing policy. If evaluating the objective function is computationally expensive, as is often the case for Graph RL problems, then using an off-policy algorithm is advisable.
@@
Symmetries that are present in the problem and solution spaces are also highly relevant and can be fruitfully exploited by algorithms and learning representations alike. For example, a TSP solution that visits nodes $(A,B,C)$ is equivalent to one that visits $(B,C,A)$. Exploiting this may lead to better sample complexity and more robust models. Approaches for canonical problems have considered encoding symmetry with specific terms in the loss function (Kwon et al., 2020; Kim et al., 2022) or augmenting the dataset of sampled solutions at inference time (Kwon et al., 2020). A recent work (Drakulic et al., 2023) proposed exploiting symmetries at the level of the MDP formulation itself by proposing a transformation of the original MDP that reduces the state space. Symmetries may also be encoded in the learning representation itself, as it can be achieved by the use of GNNs with appropriate permutation-invariant readout functions (Khalil et al., 2017).
5.1.4 Scalability to Large Problem Instances
@@
Let us enumerate some further possibilities beyond those outlined in the challenges above. One can study the application of models trained on small problem instances to larger ones. This technique requires models that are independent of the size of the problem, and can obtain impressive results in scenarios in which the model transfers well, but can equally suffer from substantial degradation in solution quality if the structure of the solution space is dissimilar at varying scales. GNNs, in particular, have proven to be an effective function approximation technique in this context, since they enable representing problem instances of different sizes transparently via subgraph embeddings (Khalil et al., 2017). Decision-time planning algorithms can be used to examine a small fraction of the entire decision-making process by focusing on constructing optimal trajectories only from the current state. Hence, instead of learning a general model for solving a wide range of instances of a graph combinatorial optimization problem, the available computational budget can be concentrated on the problem instance at hand (Darvariu et al., 2023a). Lastly, another approach for improving scalability involves using demonstrations of a well-performing algorithm to collect data instead of online environment interaction (Bhardwaj et al., 2017). This means that the expensive trial-and-error process associated with RL can be partially circumvented, and further training and fine-tuning can be performed starting from near-optimal trajectories.
5.1.5 Generalization to Unseen Problem Instances
An important issue is the fact that one cannot guarantee that the learned models will generalize well when encountering instances outside of the training distribution. This fundamental limitation is reminiscent of the No Free Lunch theorems (Wolpert & Macready, 1997) in Supervised Learning, which suggest that it is not possible to obtain a model that performs well in all possible scenarios. Partial mitigation may be possible by training models on a wide variety of scenarios including different initial starting points for the RL agent in the environment (i.e., different initial solutions), problem instances of varying difficulty and size, and even different variants of related problems.
Many works in Graph RL diversify training conditions as a means of obtaining models that generalize. This strategy is often successful (Khalil et al., 2017) but can fail in unexpected ways; for example, Darvariu et al. (2021a) found that the same RL approach for graph construction generalizes well to unseen larger instances using one objective function (resilience of the graph to random node failures), but performance collapses using a closely related objective (resilience to targeted node attacks).
To the best of our knowledge, there are currently no methods able to guarantee robustness to unexpected distribution shifts or adversarial perturbations, even for restricted classes of graph combinatorial optimization problems. In settings where such models are deployed, one may want to have the tools in place to detect distribution shift, and possibly use a fall-back approach whose properties and expected performance are well-understood.
@@ 5.2 Applications of Graph RL
6 Practical Considerations
Starting from the analysis of the existing literature and current research trends, we now derive a series of practical considerations about when and why to use (and not to use) RL for combinational optimization problems over graphs.
6.1 When and Why to Use Graph RL
When are RL approaches useful, and why might they outperform non-RL methods? Three high-level reasons that do not depend on the specific characteristics of the specific problem at hand are given below.
Flexibility regarding objective function: RL methods place few requirements on the reward function and therefore, in the context of combinatorial optimization problems, on the objective function to be optimized. RL can therefore accommodate objectives that are not “well-behaved” mathematically speaking such as non-differentiable, non-convex, and non-linear functions. It does not even require the objective function to be expressed analytically as long as samples can be generated. Exact methods, on the other hand, cannot be applied if the objective function does not belong to a pre-specified function class. Therefore, for such problems, heuristics or metaheuristics need to be used, compared to which RL can perform better due to the reasons outlined below.
Longer decision horizon: metaheuristics (e.g., greedy search, simulated annealing, evolutionary algorithms) typically use a shallow decision horizon. They move through the search space by small, local modifications of the solution. This means that, if a desirable component of the solution is unlikely to be reached by a sequence of locally optimal modifications, it may be challenging for such methods to discover them. In contrast, RL explicitly models the impact of longer sequences of actions. It is able to learn policies that, while they may not lead to large returns in the short term, will typically lead to larger expected returns over a longer decision horizon. As a relevant example, consider a network design scenario in which one aims to produce a graph with small average shortest path lengths. The ideal topology in this case is a star, but the path lengths are undefined until the network becomes connected. Methods with short decision horizons therefore struggle, while RL does not. This is indeed the case for the network design problems covered in Section 3.2, for which RL proves superior over a host of metaheuristics.
Training stage as problem-specific tuning: Another aspect that enables RL methods to outperform metaheuristics is the fact that they undergo a training stage. At a high level, this may be seen as tuning the parameters of a meta-algorithm on the particular distribution dictated by the search space of a given problem. Therefore, after training, an RL model already contains knowledge about the problem that can be used in the process of constructing a solution for an instance that, while not identical to those it has seen during training, comes from the same distribution on which the model was fit. In contrast, metaheuristics start from scratch. In practice, after model training, this translates to solutions of the same quality being found more efficiently by RL compared to generic methods, or better solutions being reached within the same computational budget.
We note that the discussion in Section 5.1.7 is also relevant but refers to the narrower goal of understanding the intricacies of the learned algorithms’ operation for an instance of a particular problem.
6.2 When and Why Not to Use Graph RL
Conversely, let us discuss situations in which RL solutions are unlikely to bring a substantial benefit over classic optimization approaches if one is interested in practical usage. For a variety of well-specified problems, especially the canonical ones, a large number of solutions and practical software tools are available. In these cases, the use of RL might be intellectually interesting, but, in practical terms, the resulting performance gain is usually very limited or absent. In general, if the objective function and decision variables of the problem are such that it can be cast into well-known paradigms, such as, for example, Integer Programming or Linear Programming, powerful and highly-optimized solvers can be leveraged.
It is also worth noting that RL does not typically yield considerable improvements if the structure of the problem is such that shallow decision horizons are sufficient for constructing optimal solutions. In this setting, modelling the expected returns of sequences of decisions, as performed by RL, is redundant.
The problem formulations considered by the current survey are such that they are amenable to RL-based approaches, which motivates the use of these methodologies in the absence of satisfactory solutions. Even in such settings, however, it is difficult to ascertain a priori how much of a gain we might obtain over a classic approach. Overall, the literature is lacking in thorough experimental comparisons between RL and non-RL methods beyond the scope of each individual paper. Furthermore, given their technical complexity and experimental challenges, RL methods may require several development cycles until satisfactory performance is obtained.
@@ 7.1 Summary
7.2 Closing Thoughts
Can we, therefore, collectively hang up our boots, leaving the machines to discover how to solve these problems? We argue that this not the case. Generic decision-making algorithms and learning representations are clearly not a silver bullet since they do not necessarily exploit problem structure efficiently. There are substantial improvements to be made by encoding knowledge and understanding about the problem into these solution approaches.
In this work, we have presented and argued for RL methodologies as an alternative to exact methods, heuristics, and metaheuristics. This dichotomy does apply when RL techniques are used in a constructive way, i.e., they build a complete solution to the problem starting from the MDP formulation as performed in the surveyed works. However, in a broader sense, RL and traditional methods are not in opposition. A number of works have explored the integration of RL and classic methods (Chen & Tian, 2019; Hottung & Tierney, 2020; Lu et al., 2020), in which the “main loop” is a traditional optimization algorithm, and RL is used for improving decisions within the method. This represents a possible path for developing algorithms that leverage both deep problem insights and highly efficient machine-learned components.
Considering the current popularity of RL, one might also ask if their application in this problem space is an instance of Maslow’s hammer (Maslow, 1966). Is it the case that we favor the ubiquitous use of this tool over careful appreciation of what might be the appropriate methodology for a given set of problems? Nevertheless, the potential of RL approaches in this problem space is significant and transformative, a belief that is strongly supported by the emerging body of literature. As RL techniques become more widespread, we expect them to find successful applications far beyond canonical problems, and to transform the scientific discovery process (Wang et al., 2023).